25 g of steam is mixed with 30g of ice, final mass of water in mixture?

Grade:12th pass

1 Answers

Arun
6 years ago

At first, 30g of ice melts down to become water. Then lets assume this 30g of water to reach steam(100°C).

L= 80 is the latent heat constant for the phase change of ice to water. C = 1 is the specific heat capacity.

For our assumption the energy released in the process is,

m(L) + mC(100-0) = 30(80) + 30(1)(100) = 5,400J

L= 540 is the latent heat constant for the phase change of steam to water.

Foe the 25g steam to condense we need a energy equal to m(L) = 25(540) = 13,500J

But in the above mentioned process only 5,400J is released which will not condense all the steam. Hence our assumption is correct.The final temperature will be 100°C.